# DAIRstega: Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel

### v4

@@ 3.1 Settings
Dataset.
We obtained a large amount of data from Wikipedia, Twitter, GPT4 (OpenAI 2023), and the publicly available data for training LLMs111 https://github.com/tloen/alpaca-lora/tree/main. The final dataset contains 108,771 texts and nearly 100 different discourse characteristics to train the language models of all schemes.
Model configuration.
DAIRstega can access any open-source LLMs. In the experiment, we used LLaMA2-7B and 13B models (Meta 2023) as an example. The rank of LoRA (Hu et al. 2021) is 8, and the dropout is 0.05. The learning algorithm is AdamW (Loshchilov and Hutter 2018), and the learning rate is 1e-5. The batch size is 16. For simplicity, we use $\beta=1,0.5$ and $\alpha=8,32,48$ to perform DAIRstega.
Table 1: The overall results of different embedding ways. The embedding rate of each way is near 1Bpw (bit per word). Bold and “ *” are the best and the suboptimal result. “LLM” in this table and Ours are all selected the LLaMA2-7B. For the LS works involved in these embedding ways, see “Section 3.1”. $\uparrow$” is that the higher the value, the better the result. $\downarrow$” means the lower the value, the better the result. The complete data can be seen in Table 8 and Table 9 in Supplementary A.2.
@@
Ours 8.030 0.342 97.43 38.60 0.227 6.341 2.57 78.53* 77.42
Embedding $\hookrightarrow$ Anti-steganalysis
Score $\uparrow$ LS_CNN TS_CSW EILG UP4LS
@@
Ours 65.41 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
Baselines.
Comparison is divided into two parts, one is the comparison between the proposed embedding way and the existing embedding ways, and the other is the comparison between DAIRstega and the prevalent works in recent years.
$\cdot$ In terms of the embedding ways, the existing LS works can be summarized as follows: (1) HC (or VLC), (2) ADG, (3) AC, and (4) FLC. (1) HC is used in VAEstega (Yang et al. 2021), LSCS (Lu et al. 2023), Hi-Stega (Wang et al. 2023a), RNNstega (Yang et al. 2019), GAN-APD (Zhou et al. 2021), JLS (Ding et al. 2023a), Discop (Ding et al. 2023b). (2) ADG is used in Hi-Stega (Wang et al. 2023a), ADG (Zhang et al. 2021). (3) AC is used in VAEstega (Yang et al. 2021), Hi-Stega (Wang et al. 2023a). (4) FLC is used in RNNstega (Yang et al. 2019), JLS (Ding et al. 2023a).
$\cdot$ In terms of the prevalent works, we selected the LS and watermarking works that showed excellent performance as the baselines. LS: (1) Fang (ACL) (Fang, Jaggi, and Argyraki 2017). (2) RNNstega (IEEE TIFS) (Yang et al. 2019). (3) ADG (ACL-IJCNLP) (Zhang et al. 2021). (4) LSCS (IJCNN) (Lu et al. 2023). Watermarking: (5) PLMmark (AAAI) (Liu et al. 2023). It uses pre-trained language models to provide the first secure black-box watermarking. They can compare the generation effects of DAIRstega from different aspects, and these performances have been widely recognized.
Evaluation metrics.
@@
$\cdot$ In terms of text quality (perceptual concealment), we adopt (1) PPL $\downarrow$ (stego perplexity). (Wu et al. 2023; Ding et al. 2023a). (2) $\Delta$Pcs $\downarrow$ (The difference between PPL of cover and stego (Yang et al. 2021)).
$\cdot$ In terms of statistical analysis (statistical concealment), we adopt (3) CS $\uparrow$. (Cosine similarity). (4) JSD $\downarrow$, which averages the KL divergence. (5) ED $\downarrow$ (Euclidean distance). (6) MD $\downarrow$ (Manhattan distance). (7) $\Delta$DP $\downarrow$ (Dot product difference (Cai et al. 2021)).
$\cdot$ In terms of discourse matching (Semantic concealment), we adopt (8) $\Delta$LDA $\downarrow$. (9) Mauve $\uparrow$ (Pillutla et al. 2021). Mauve is a metric that considers fluency and consistency. (10) BLEU $\uparrow$ (Bilingual evaluation understudy (Papineni et al. 2002)). (11) Score (BERTScore) $\uparrow$ (Zhang et al. 2020). Score considers the similarity at the segment level.
$\cdot$ In terms of anti-steganalysis, we use the high-performance linguistic steganalysis LS_CNN (Wen et al. 2019), TS_CSW (Yang, Huang, and Zhang 2020), EILG (Xu, Zhang, and Liu 2023), and UP4LS (Wang et al. 2023b): (12) Acc $\downarrow$ (Accuracy) and (13) F1 $\downarrow$ (F1-score).
3.2 Comparison of the embedding ways
To illustrate the effect of embedding way, we performed this experiment. We add the existing works with the same LLMs as ours. The results are shown in Table 1.
The results in Table 1 show that the proposed embedding way can improve the stego quality, better simulate the CP distribution of covers, and get the optimal anti-steganalysis. Overall performance performs better than the existing ways.
3.3 Comparison with LS baselines
Text quality (Perceptual concealment).
Table 2 shows the comparison regarding the stego quality.
Table 2: Text quality comparison of stegos generated by DAIRstega and LS baselines. Bold is the best result. “ *” is the suboptimal result. $\alpha$ and $\beta$ can be found in Formula 9. The meanings of the parameters “bin”, “bit”, and “$\tau$” are the same in Fang, RNNstega, and ADG schemes.
@@ 3.3 Comparison with LS baselines
Statistical analysis (Statistical concealment).
@@
Table 3: Statistical analysis comparison of stegos generated by DAIRstega and LS baselines. The meanings of “Bold”, “ *”, “bin”, “bit”, and “$\tau$” are the same in Table 2.
@@
Discourse matching (Semantic concealment).
Table 4 shows the comparison regarding discourse matching.
Table 4: Discourse-matching comparison of stegos generated by DAIRstega and LS baselines. Unit is %. The complete data can be seen as Table 11 in Supplementary A.3.
@@
Table 5: Anti-steganalysis (Acc $\downarrow$) comparison of stegos generated by DAIRstega and LS baselines. Unit is %. The complete data can be seen as Table 12 in Supplementary A.3. We give the detection of the other 7 steganalysis (including non-BERT, BERT, and LLMs-based). It can be seen as Table A.4 in Supplementary A.4.
Schemes Param LS_CNN TS_CSW EILG UP4LS
Fang bin = 1 99.70 88.34 99.50 99.75
@@
3.4 Comparison with watermarking baseline
The rise of LLMs has spawned many excellent watermarking papers (Liu et al. 2023; Kirchenbauer et al. 2023). They provide solutions for copyright protection in terms of high fidelity (Kirchenbauer et al. 2023), consistency (Zhang et al. 2024), and abuse resistance (Liu et al. 2023). Excitingly, DAIRstega can also be regarded as an invisible watermarking with strong fidelity, consistency, and abuse resistance.
The focus of this paper is not watermarking, we only compare it with the PLMmark (AAAI 2023), as shown in Table 6.
Table 6: Fidelity, consistency, and abuse resistance comparison with watermarking baseline. The complete data can be seen in Table 11 and Table 12 in Supplementary A.3.
[…]
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$
PLMmark 18.748 11.056 0.009 7.22
DAIRstega 7.686 0.006 0.006 80.83
Scheme $\hookrightarrow$ Abuse resistance
Score $\uparrow$ TS_CSW $\downarrow$ EILG $\downarrow$ UP4LS $\downarrow$
PLMmark 43.50 98.81 99.92 99.85
DAIRstega 65.76 54.05 62.28 69.04
According to the comparison results in Table 6, the watermark texts generated by DAIR have excellent fidelity and consistency, and it is difficult for unauthorized receivers to detect the existence of the watermark text, avoiding the problem of watermark abuse. Only authorized receivers with the same model settings can detect and fully extract the watermark information.
@@ 3.5 Longer stegos by DAIRstega
Table 7: Examples of long stegos generated by DAIRstega. Bitstream is the binary of the secret. Secret can be not only texts, but also other modalities. In addition, we embed the “AAAI 2025 icon”. For details, see Table 14 in Supplementary A.5.
@@
The secret will change the next token. This makes it difficult to maintain logical coherence when generating longer stegos. So longer stegos with high quality is a problem, and there is little research in this task. Fortunately, we found that DAIRstega also has this ability. Examples of embedded text and cross-modal secrets are shown in Table 7 and Table 14.
3.6 Ablation experiment
We performed ablation experiment on different LLMs, and the results are shown in Table 8.
@@ 3.6 Ablation experiment
$\beta$=1, $\alpha$=32 PPL $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$
LLaMA2-7B 8.184 97.65 38.48 0.217
LLaMA2-13B 8.431 96.32 43.93 0.271
$\beta$=1, $\alpha$=32 MD $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$
LLaMA2-7B 6.293 0.022 76.97 77.85
LLaMA2-13B 8.034 0.019 88.38 72.31
$\beta$=1, $\alpha$=32 $\Delta$DP $\downarrow$ Score $\uparrow$ TS_CSW $\downarrow$ UP4LS $\downarrow$
LLaMA2-7B 2.35 65.22 54.05 71.97
LLaMA2-13B 3.68 61.66 54.73 71.12
According to Table 8, it is not possible to generate stegos with stronger concealment by using larger LLMs. This is because LS needs to embed secret. The stego quality is more relevant with embedding way rather than using larger LLMs.
@@ 4 Conclusion
To enhance the concealment of stegos, this paper proposes an LS scheme based on dynamically allocated intervals, namely DAIRstega. It uses the idea of non-uniform roulette and takes the CP of tokens as an important basis for allocating the number of codings. Then the requirements of invisible information hiding are analyzed and three constraints are given. The allocation expression is condensed. Experiments show that the proposed embedding way and DAIRstega have achieved excellent performance in various concealment and anti-steganalysis. We also prove that the stego quality depends more on the design of the model and embedding way, not just the use of LLMs. Furthermore, DAIRstega can be used as a secure watermarking scheme. Last but not least, to concisely show the proposed scheme, we did not conduct the results of all functions and LLMs.

### v5

@@ 4.1 Settings
Dataset
We obtained a large amount of data from Wikipedia, Twitter, GPT4 [18], and the publicly available data for fine-tuning LLMs111 https://github.com/tloen/alpaca-lora/tree/main. The final dataset contains 57,414 texts and nearly 100 different discourse characteristics to fine-tune the language models of all schemes.
MindSpore framework
All experiments are based on the MindSpore framework. MindSpore is an open-source AI framework introduced by Huawei, designed to support flexible deployment across cloud, edge, and device scenarios. In this research, MindSpore played a significant role in unleashing the computational power of the hardware.
Model configuration
LLM is divided into closed-source and open-source LLMs, closed source such as Gemini, GPT4, etc., open-source such as LLaMA2-7B, etc. For detailed information on closed-source LLMs, please refer to the review of comprehensive research in related fields [19, 20]. Since DAIRstega needs to modify the token sampling inside LLMs, this paper uses open-source LLMs for experiments. DAIRstega can access any open-source LLMs. In the comparison experiments, the LLaMA2-7B model is used [21] as the base LLM. The LLaMA2-7B and LLaMA2-13B models are used in the ablation experiments of the different LLMs, and all models are operated with 16-bit precision. For simplicity, we use $\beta=1,0.5$ and $\alpha=8,32,48$ to perform DAIRstega.
Baselines
Since the steganography scheme consists of a language model and an embedding way, the language models of existing steganography are different. Therefore, to ensure fairness in the comparison, we not only compared the schemes, but also compared the performance of different embedding ways. That is to say, comparison is divided into two parts: one is the comparison between the proposed embedding way and the existing embedding ways, and the other is the comparison between DAIRstega and the prevalent works in recent years.
Table 1: Overview of each part of the experiments. It includes: experimental part (Focus), Section, Metrics, Base LLM, Table, and supplementary materials for table (Supplement).
Focus Section Metrics Base LLM Table Supplement
Comparison study Embedding ways 4.2 Perceptual, Statistical, Semantic, Anti-steganalysis LLaMA2-7B 2 10, 11
Steganography baselines 4.3.1 Perceptual 3
4.3.2 Statistical 4
4.3.3 Semantic 5 12
4.3.4 Anti-steganalysis 6 13, 7.3
Related-task baseline 4.4 Quality, Consistency, Unforgeability 7
DAIRstega exploration Longer stegos 4.5 Perceptual, Semantic LLaMA2-7B 8 15
Ablation of LLMs 4.6 Perceptual, Statistical, Semantic, Anti-steganalysis LLaMA2-7B, 13B 9
Table 2: The overall results of different embedding ways (based on MindSpore framework). The embedding rate of each way is near 1Bpw (bit per word). Bold and “ *” are the best and the suboptimal result. The “LLM”s in this table are all LLaMA2-7B. For the steganography works involved in these embedding ways, see “Section 4.1”. $\uparrow$” is that the higher the value, the better the result. $\downarrow$” means the lower the value, the better the result. The complete data can be seen in Table 10 and Table 11 in Appendix 7.1.
@@ 4.3.3 Semantic 5 12
LLM+Ours 8.030 0.342 97.43 38.60 0.227 6.341 2.57 78.53* 77.42
Embedding ways $\hookrightarrow$ Anti-steganalysis
Score $\uparrow$ LS_CNN [30] TS_CSW [31] EILG [32] UP4LS [33]
@@
LLM+Ours 65.41 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
  • 1. HC (Huffman coding) is used in VAE-Stega [3], LSCS [4], RNNstega [9], GAN-APD [11], JLS [10], and Discop [17].
  • 2. ADG (Adaptive dynamic grouping) is used in Hi-Stega [5] and ADG [12].
  • 3. AC (Arithmetic coding) is used in VAE-Stega [3], Hi-Stega [5], and Meteor [13].
  • 4. FLC (Fixed length coding) is used in RNNstega [9] and JLS [10].
$\cdot$ In terms of the embedding ways, the existing works can be summarized as follows: (1) Huffman coding (HC) (or Variable length coding (VLC)), (2) Adaptive dynamic grouping (ADG), (3) Arithmetic coding (AC), and (4) Fixed length coding (FLC). (1) HC is used in VAE-Stega [3], LSCS [4], RNNstega [9], GAN-APD [11], JLS [10], Discop [17]. (2) ADG is used in Hi-Stega [5] and ADG [12]. (3) AC is used in VAE-Stega [3], Hi-Stega [5], and Meteor [13]. (4) FLC is used in RNNstega [9], JLS [10].
$\cdot$ In terms of the prevalent works, we selected the steganography and watermarking works that showed excellent performance as the baselines. Steganography-task baselines: (1) Fang [16]. (2) RNNstega [9]. (3) ADG [12]. (4) LSCS [4]. Related-task baseline: (5) PLMmark [22]. It uses pre-trained language models to provide the first secure black-box watermarking scheme. They can compare the generation effects of DAIRstega from different aspects, and these performances have been widely recognized.
Evaluation metrics
@@
$\cdot$ In terms of perceptual concealment (text quality), we adopt (1) PPL $\downarrow$ (stego perplexity). [23, 10]. (2) $\Delta$Pcs $\downarrow$ (The difference between PPL of cover and stego [3]).
$\cdot$ In terms of statistical concealment (statistical analysis), we adopt (3) CS $\uparrow$. (Cosine similarity). (4) JSD $\downarrow$, which averages the KL divergence. (5) ED $\downarrow$ (Euclidean distance). (6) MD $\downarrow$ (Manhattan distance). (7) $\Delta$DP $\downarrow$ (Dot product difference [24]).
$\cdot$ In terms of semantic concealment (discourse matching), we adopt (8) $\Delta$LDA $\downarrow$. (9) Mauve $\uparrow$ [25]. Mauve is a metric that considers fluency and consistency. (10) BLEU $\uparrow$ (Bilingual evaluation understudy [26]). (11) Score (BERTScore) $\uparrow$ [27]. Score considers the similarity at the segment level. $\cdot$
In terms of anti-steganalysis, steganalysis is essentially a detection method [28]. It can detect whether there is secret information in the text [29]. we use the high-performance linguistic steganalysis LS_CNN [30], TS_CSW [31], EILG [32], and UP4LS [33]: (12) Acc $\downarrow$ (Accuracy) and (13) F1 $\downarrow$ (F1-score).
Experimental overview
We provide an overview of each part of the experiments, as shown in Table 1.
4.2 Comparison with the different embedding ways
To illustrate the effect of the embedding way, we performed this experiment. We add the existing works with the same LLMs as ours. The results are shown in Table 2.
Table 2 shows that the proposed embedding way can improve the stego quality (perceptual concealment), better simulate the CP distribution of covers (statistical concealment), and get the optimal anti-steganalysis. Overall performance performs better than the existing ways.
4.3 Comparison with steganography-task baselines
4.3.1 Perceptual concealment (text quality)
Table 3 shows the comparison regarding the perceptual concealment.
Table 3: Perceptual concealment comparison of stegos generated by DAIRstega and steganography baselines. Bold is the best result. “ *” is the suboptimal result. $\alpha$ and $\beta$ can be found in Formula 9. The meanings of the parameters “bin”, “bit”, and “$\tau$” are the same in Fang, RNNstega, and ADG schemes.
@@ 4.3.1 Perceptual concealment (text quality)
4.3.2 Statistical concealment (statistical analysis)
@@
Table 4: Statistical concealment comparison of stegos generated by DAIRstega and steganography baselines. The meanings of “Bold”, “ *”, “bin”, “bit”, and “$\tau$” are the same in Table 3.
@@
4.3.3 Semantic concealment (discourse matching)
Table 5 shows the comparison regarding semantic concealment.
Table 5: Semantic concealment comparison of stegos generated by DAIRstega and steganography baselines. The unit is %. The complete data can be seen as Table 12 in Appendix 7.2.
@@ 4.3.4 Anti-steganalysis
Table 6: Anti-steganalysis (Acc $\downarrow$) comparison of stegos generated by DAIRstega and steganography baselines. The unit is %. The complete data can be seen as Table 13 in Appendix 7.2. We also give the detection of the other 7 steganalysis works [14, 34, 35, 36, 37, 38, 15] (including non-BERT, BERT, and LLM-based). It can be seen as Table 7.3 in Appendix 7.3.
Schemes Param LS_CNN [30] TS_CSW [31] EILG [32] UP4LS [33]
Fang [16] bin = 1 99.70 88.34 99.50 99.75
@@
4.4 Comparison with related-task baseline
The rise of LLMs has spawned many excellent watermarking papers [22, 39]. They provide solutions for copyright protection in terms of high text quality [39], consistency [40], and unforgeability [22]. Excitingly, DAIRstega can also be regarded as an invisible watermarking with high quality, consistency, and unforgeability.
The focus of this paper is not watermarking, we only compare it with the PLMmark, as shown in Table 7.
Table 7: Quality, consistency, and unforgeability comparison with watermarking baseline. The complete data can be seen in Table 12 and Table 13 in Appendix 7.2.
Scheme Quality Consistency Unforgeability
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ Score $\uparrow$ TS_CSW $\downarrow$ EILG $\downarrow$ UP4LS $\downarrow$
PLMmark [22] 18.748 11.056 0.009 7.22 43.50 98.81 99.92 99.85
DAIRstega $\beta$=1, $\alpha$=8 7.686 0.006 0.017* 79.91* 65.76 56.82 67.25 75.40
$\beta$=1, $\alpha$=32 8.184* 0.492* 0.022 76.97 65.22 54.05 62.28 71.97
$\beta$=1, $\alpha$=48 8.219 0.527 0.020 78.70 65.25* 54.98* 65.51* 71.40*
$\beta$=0.5, $\alpha$=8 13.775 6.082 0.018 80.83 62.49 63.13 65.67 69.04
$\beta$=0.5, $\alpha$=32 20.876 13.184 0.029 75.58 60.89 68.73 71.71 76.41
$\beta$=0.5, $\alpha$=48 22.069 14.377 0.006 72.40 60.29 69.48 70.64 79.46
According to the comparison results in Table 7, the watermark texts generated by DAIR have excellent quality and consistency, and it is difficult for unauthorized receivers to detect the existence of the watermark text, avoiding the problem of watermark abuse. Only authorized receivers with the same model settings can detect and fully extract the watermark information.
@@ 4.5 Longer stegos by DAIRstega
Table 8: Examples of long stegos generated by DAIRstega. Bitstream is the binary of the secret. Secret can be not only texts, but also other modalities. In addition, we embed the image information. For details, see Table 15 in Appendix 7.4.
@@
The secret will change the next token. This makes it difficult to maintain logical coherence when generating longer stegos. So, longer stegos with high quality is a problem, and there is little research in this task. Fortunately, we found that DAIRstega also has this ability. Examples of embedded text and cross-modal secrets are shown in Table 8 and Table 15.
4.6 Ablation of the different LLMs
We performed ablation experiments on different LLMs, and the results are shown in Table 9.
@@ 4.6 Ablation of the different LLMs
$\beta$=1, $\alpha$=32 Perceptual Statistical
PPL $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$
LLaMA2-7B 8.184 97.65 38.48 0.217 6.293 2.35
LLaMA2-13B 8.431 96.32 43.93 0.271 8.034 3.68
$\beta$=1, $\alpha$=32 Semantic Anti-steganalysis
Mauve $\uparrow$ BLEU $\uparrow$ $\Delta$LDA $\downarrow$ Score $\uparrow$ TS_CSW $\downarrow$ UP4LS $\downarrow$
LLaMA2-7B 76.97 77.85 0.022 65.22 54.05 71.97
LLaMA2-13B 88.38 72.31 0.019 61.66 54.73 71.12
According to Table 9, it is not possible to generate stegos with stronger concealment by using larger LLMs. This is because steganography needs to embed secrets. The stego quality is more relevant with the embedding way rather than using larger LLMs.
@@ 5 Conclusion
To enhance the concealment of stegos, this paper proposes a steganography scheme based on allocated roulette area, namely DAIRstega. It uses the idea of the non-uniform roulette wheel and takes the CP of tokens as an important basis for allocating the number of codings. Then, the allocation function and three constraints are designed to optimize the allocation process. Experiments show that the proposed embedding way and DAIRstega have achieved excellent performance in various concealment and anti-steganalysis. Last but not least, to show the proposed scheme concisely, we did not conduct the results of all allocation functions and LLMs.